OpenStudy (anonymous):

verify sint/1-cost=csct+cott so far I got the RHS to equal 1+cost/sint how can I make both sides equal without crossing over?

OpenStudy (anonymous):

you're basically done.. :)

OpenStudy (anonymous):

\(\large \frac{1+cost}{sint} =\frac{1}{sint} + \frac{cost}{sint} = \) ????

OpenStudy (anonymous):

\[\frac{ 1 }{ sint }= csct \]

OpenStudy (anonymous):

than

OpenStudy (anonymous):

\[\frac{ cost }{ sint }=cott\]

OpenStudy (anonymous):

well in my case, when i am stuck on RHS, i see LHS and see what i need whether i can find some connection

OpenStudy (anonymous):

what do you think you'll need to do in order to make \[(1+cosx)/sinx\] into\[sinx/(1-cosx)\]

OpenStudy (anonymous):

what equation has both sinx and cosx in it?

OpenStudy (anonymous):

\[\sin ^{2}t+\cos ^{2}t=1\]

OpenStudy (anonymous):

well even i don't know how i got to this.. i'll just show you\[(1+cosx)/sinx = sinx(1+cosx)/\sin^2x = sinx(1+cosx)/(1-\cos^2x)\]

OpenStudy (anonymous):

\[\frac{ sint }{ cost }=tant\]

OpenStudy (anonymous):

you know how to do the rest right?

OpenStudy (anonymous):

let me right it out

OpenStudy (anonymous):

what kind of book has these difficult trigonometry questions?

OpenStudy (anonymous):

swokowski algebra and trig

OpenStudy (anonymous):

all these questions could be on our test tomorrow

OpenStudy (anonymous):

i see. good luck:)

OpenStudy (anonymous):

thx i need it

OpenStudy (anonymous):

so you multiplied the right side with sinx/sinx?

OpenStudy (anonymous):

sorry but im stuck, not sure how you got to where you got. where did the sinx come from?

OpenStudy (anonymous):

\[\frac{\sin t}{1-\cos t} = \csc t+\cot t\] Because csc t = 1 / sin t, and cot t = cos t / sin t this becomes: \[\frac{\sin t}{1-\cos t} = \frac{1+\cos t}{\sin t}\] Multipling left side by sin t/sin t and right side by (1- cos t)/(1- cos t) \[\frac{\sin^{2}t}{\sin t\left(1-\cos t\right)} = \frac{\left(1+\cos t\right)\left(1-\cos t\right)}{\sin t\left(1-\cos t\right)}\] Expanding brackets gives \[\frac{\sin^2{t}}{\sin t\left(1-\cos t\right)} = \frac{1-\cos^{2}t}{\sin t\left(1-\cos t\right)}\] We know that sin^2 t + cos^2 t = 1, so sin^2t = 1 - cos^2t: \[\frac{1-\cos^{2}t}{\sin t\left(1-\cos t\right)} = \frac{1-\cos^{2}t}{\sin t\left(1-\cos t\right)}\]

OpenStudy (anonymous):

thanks a million!

OpenStudy (anonymous):

sorry i'm late. Meepi is right. again sinx/sinx =1, an identity.

OpenStudy (anonymous):

thanks